OpenStudy (ltrout):

[MEDAL] Find the maximum value of the function for the polygonal convex set determined by the given system of inequalities

OpenStudy (ltrout):

Equations are as follows: \[3x-2y \ge 0\] \[y \ge 0\] \[3x+2y \le 24\]

OpenStudy (ltrout):

Vertices of the triangle formed by those equations are: (0, 0), (4,6), (8,0)

OpenStudy (ltrout):

System of inequalities is: \[f(x,y)=7y-3x\]

OpenStudy (instagrammodel):

Oh God! Its these Lol.

OpenStudy (ltrout):

@InstagramModel You know how to solve?

OpenStudy (instagrammodel):

it's been awhile. But.. Lemme think.

OpenStudy (ltrout):

I know I have to plug in one of those vertices for the f(x,y) equation, but I'm not sure which.

OpenStudy (instagrammodel):

im sorry. I honestly cannot remember how to do this :( im sorry